OpenStudy (anonymous):

(x+1)/(x)=(-7)/(x-12)

OpenStudy (anonymous):

(x+1)/(x)=(-7)/(x-12) (x+1)(x-12)=-7x x^2-11x-12=-7x x^2-4x-12=0 (x-6)(x+2)=0 x=6,-2

OpenStudy (anonymous):

he is right

OpenStudy (anonymous):

thanks this was on my practice algebra regents